OpenStudy (wlhfwqfhow):

A student writes an incorrect step while checking if the sum of the measures of the two remote interior angles of triangle ABC below is equal to the measure of the exterior angle. Step 1: m∠a + m∠b + m∠c = 180° (sum of angles of a triangle) Step 2: m∠d − m∠c = 90° (complementary angles) Step 3: Therefore, m∠a + m∠b + m∠c = m∠d + m∠c Step 4: So, m∠a + m∠b = m∠d

OpenStudy (will.h):

mostly in step 2 it would be m<d + m<c = 90

OpenStudy (will.h):

if it is a straight line then m<d + m<c = 180

OpenStudy (will.h):

there should be a diagram or anything like that tho

OpenStudy (welshfella):

yea it would make it easier.

OpenStudy (welshfella):

I guess <d must be the exterior angle,

OpenStudy (will.h):

here's what i think it should be Step 1: m∠a + m∠b + m∠c = 180° (sum of angles of a triangle) Step 2: m∠d + m∠c = 180° (Supplementary angles) Step 3: Therefore, m∠a + m∠b + m∠c = m∠d + m∠c (because 180 = 180) Step 4: So, m∠a + m∠b = m∠d